OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3.

OpenStudy (anonymous):

Let g be f inverse. If f(x) = 3, then g(f(x)) = g(3), which is what we want to find. Since g is the inverse, g(f(x)) = x. So, we want x such that f(x) = 3. Then, 3 = 4(x+3) - 5. 8 = 4x +12, -4 = 4x, -1 = x

OpenStudy (anonymous):

Ah! You're such a great help. Thank you so so much!!